# Using VLM Reasoning to Constrain Task and Motion Planning

### v1

@@ Abstract
In task and motion planning, high-level task planning is done over an abstraction of the world to enable efficient search in long-horizon robotics problems. However, the feasibility of these task-level plans relies on the downward refinability of the abstraction into continuous motion. When a domain’s refinability is poor, task-level plans that appear valid may ultimately fail during motion planning, requiring replanning and resulting in slower overall performance. Prior works mitigate this by encoding refinement issues as constraints to prune infeasible task plans. However, these approaches only add constraints upon refinement failure, expending significant search effort on infeasible branches. We propose VIZ-COAST, a method of leveraging the common-sense spatial reasoning of large pretrained Vision-Language Models to identify issues with downward refinement a priori, bypassing the need to fix these failures during planning. Experiments on two challenging TAMP domains show that our approach is able to extract plausible constraints from images and domain descriptions, drastically reducing planning times and, in some cases, eliminating downward refinement failures altogether, generalizing to a diverse range of instances from the broader domain.
@@ I Introduction
Task and Motion Planning (TAMP) [1] offers a framework for solving long-horizon robotic problems that couples symbolic decision-making with geometric feasibility, such as clearing a cluttered table or navigating between rooms where a key must first be found to open a door. These problems are compelling not only because completing tasks requires multiple steps, but also because they demand reasoning across different levels of abstraction. The robot must determine what actions to take and how to execute them. This feature makes TAMP inherently challenging as it requires reconciling discrete task structures with high-dimensional geometric constraints.
@@
The two levels of abstraction, task and motion, enable efficient planning in long-horizon problems. At the task level, actions are represented symbolically, such as “pick,” “place,” or “navigate,” with each action specified by preconditions and effects on a set of symbolic predicates. This abstraction allows the planner to efficiently reason about possible sequences of actions without searching in continuous geometric space. However, the core difficulty is that these abstractions often cannot express critical geometric information, violating the downward refinement property [2, 3]: a valid task plan may not necessarily correspond to a sequence of feasible motion plans. For instance, a symbolic plan may prescribe grasping an object, but that object may be obstructed by another object, preventing any collision-free grasp.
@@
Figure 1: VIZ-COAST uses VLMs and SMT-based task planning to constrain TAMP problems. Our constraints generalize to unseen problem instances within a broader domain zero-shot. VIZ-COAST introduces minimal search overhead, enabling more efficient planning than prior state-of-the-art.
@@
PDDLStream [4] is a TAMP framework which circumvents downward refinement failures by employing a sample-first approach, in which feasible continuous action parameterizations, such as grasp poses and trajectories, are sampled prior to planning via streams, allowing task planning to be done with geometric feasibility pre-certified. However, PDDLStream expends significant effort sampling parameterizations of action groundings which may not be used in the final plan. This can become highly inefficient, as the number of possible action groundings grows exponentially with the number of objects.
@@
Alternatively, many TAMP methods take a plan-first approach, creating a complete task plan before attempting downward refinement. Intuitively, it is far more effective to prune the search space at the discrete task level than the continuous motion level, especially in long-horizon tasks. When refinement fails, plan-first methods translate motion failures to task-level constraints before replanning. For example, upon failing to find a feasible grasp pose due to an obstructing object, a constraint may be added to enforce that the offending object must not be in that position when attempting a grasp. IDTMP [5] implements this by formulating planning as with SATPlan [6, 7] through Z3 [8], a Satisfiability Modulo Theories (SMT) solver, incrementally adding constraint clauses upon failure. COAST [9], a plan-first successor to PDDLStream, adds auxiliary predicates and entities to directly encode constraints in the PDDL description. All these existing constraint-based approaches ultimately require expending additional search effort to encounter downward-refinement failures and discover the necessary constraints, often replanning several times before finding a plan that can be successfully ground to motion. Our goal is to avoid expending this effort by adding the necessary constraints before planning even begins.
@@
Recently, large pretrained vision language models (VLMs) have demonstrated common-sense geometric and physical reasoning capabilities [10, 11], allowing them to generalize zero-shot to a wide range of settings. For this reason, they have been employed for various facets of TAMP, including subgoal generation [12], encoding open-world preferences [13], and correcting errors [14].
@@
In this work, we propose Visual Insight for Z3 [8]-based Constraints and Streams (VIZ-COAST), a novel approach to TAMP which leverages VLMs to foresee potential downward-refinement failures and add constraints to avoid them before planning begins, eliminating the need to encounter these failures during planning and expend effort replanning.
@@
Through experiments on two challenging TAMP domains, we demonstrate that VIZ-COAST significantly decreases planning time by reducing or even eliminating replanning caused by downward refinement failures. We also show that our constraints consistently generalize zero-shot to unseen problem instances within a broader domain.
@@ II Background and Related Work
Figure 2: The VIZ-COAST architecture. Our Visual Reasoning Module takes as input an example scene, consisting of an image and the geometric state, the PDDL domain description, and an example of a syntactically compliant output. It produces a Python file encoding constraints through a high-level API to the SMT-based task planner. When presented with a new problem instance, the task planner takes these constraints as input, in addition to the PDDL domain and new problem descriptions. The planner applies the constraints to block geometrically infeasible task plans. The high-level plan produced is then grounded by a streams-based motion planner to produce an executable continuous motion plan.
In this section, we begin with a general overview of modern task and motion planning research and its main challenges in Sec.˜II-A. We then review prior approaches in Sec.˜II-B that accelerate TAMP by constraining the search space, highlighting the need for a generalizable method capable of pruning before planning begins. Lastly, in Sec.˜II-C, we explain how VLMs’ spatial reasoning capabilities make such a method possible and clarify how our approach differs from prior work.
@@ II-A Task and Motion Planning
Task and motion planning (TAMP) decomposes continuous long-horizon motion planning into search in both discrete (task-level) and continuous (motion-level) spaces [1]. The task level planning is done over an abstraction of the robot’s capabilities, and produces a sequence of symbolic actions to transition from a given start state to a desired goal condition (e.g., STRIPS [15]). In this work we formalize task planning using Planning Domain Definition Language (PDDL) [16] due to its expressive power and widespread adoption.
A core challenge in TAMP is downward-refinement failure [2, 3], which occurs when a symbolic task plan becomes infeasible at the motion level. PDDLStream [4] addresses this issue by taking a sample-first approach, performing continuous sampling before task planning via declarative streams, allowing earlier reasoning about geometric feasibility. Other works [17, 18, 19, 9] take a plan-first approach, designing algorithms to add constraints upon encountering failures which inform task-level replanning to prevent revisiting infeasible options. In the next section, we expand on the role of constraints and other methods of reducing the search space in TAMP.
@@ II-B Constraints in TAMP
The purpose of constraints in both IDTMP and COAST is to prune the task-level search space to avoid repeating failures, though they differ in implementation. COAST [9] is a plan-first adaptation of PDDLStream which encodes constraints by modifying the PDDL description, adding auxiliary predicates and entities to block actions under specific conditions. While this approach allows COAST to be planner-agnostic, encoding constraints directly in PDDL limits the constraints that may be expressed, and the auxiliary constructions introduce significant search overhead. PDDL 3.0 [20] introduced constraints on valid states, but is poorly supported by existing planners (e.g., Fast Downward [21]) and cannot be reliably generated by LLMs due to limited training data. In contrast, we take the approach done by IDTMP [5], and solve for valid task plans through SATPlan [6, 7] encoded in a Satisfiability Modulo Theories (SMT) solver (in particular, Z3 [8]). This formulation enables constraints to be directly encoded in the SMT solver through simple Python programming, enabling efficient and flexible constraint specification. While heuristic search approaches typically exhibit faster planning times [22], SATPlan encodings offer far greater specification flexibility and can avoid the costly translation overhead in FastDownward [21] for many problems.
Although existing methods [9, 5] are able to overcome downward refinement issues, they discover constraints only by encountering motion grounding failures during execution. This reactive process is often repeated for several iterations before uncovering the necessary constraints to solve a particular problem instance, and cannot generalize to unseen instances.
Alternative approaches have explored predicting action feasibility to avoid refinement failures. Wells et al. [23] learn feasibility constraints for tabletop manipulation through domain-specific training, while Xu et al. [24] employ neural feasibility checks to identify infeasible actions before motion planning. Driess et al. [25] use deep visual heuristics to assess feasibility in mixed-integer manipulation planning, and [26] leverage visual reasoning to predict valid actions for TAMP. Although promising, these methods require extensive domain-specific training data and may not generalize well to new environments or object configurations. In contrast, recent advances in LLMs and VLMs enable the use of pre-trained common-sense priors to reduce unnecessary exploration.
@@ II-C Vision Language Models for TAMP
LLMs have demonstrated that they are suitable for high [27]. VLMs further integrate visual and textual inputs to enable scene understanding [28], spatial reasoning [11], and grounded action execution [29, 14]. These capabilities make VLMs especially promising for addressing the geometric reasoning challenges in TAMP.
Several works have integrated LLMs and VLMs into TAMP pipelines for various purposes. Zhang et al. [30] use VLMs to ground task planners by validating symbolic predicates and providing incremental feedback. Hu et al. [31] uses a VLM to generate physically feasible actions. ReplanVLM [32] employs VLMs for plan generation with replanning upon grounding failures. CaStL [33] translates natural language task descriptions into PDDL with constraints reflecting user preferences. OWL-TAMP [13] encodes open-world preferences as constraints on symbolic and geometric states. Both CaStL and OWL-TAMP operate at the problem specification level, aiming to capture preferences rather than prevent downward refinement failures. VLM-TAMP [12] uses VLMs’ common-sense knowledge to generate intermediate subgoals for long-horizon planning; however, it assumes plans can always be refined at the motion level, making it prone to failure when this assumption fails. CLIMB [34] incrementally updates the domain description upon failure with an LLM to continually improve task performance; however, it is restricted to constraints expressible in PDDL, and does not make use of visual input.
@@ III Problem Formulation
Our work investigates the use of VLMs to accelerate solving the TAMP problem. In particular, our method is capable of producing constraints which generalize to a range of problem instances within a broader domain. When inferring constraints for a particular domain, we assume access to a model of the robot and its environment, as well as a Planning Domain Definition Language (PDDL) [35] domain description which specifies the available actions, predicates, and possible objects. Additionally, we assume access to an image of some example scene within the domain. We assume knowledge of the underlying geometric state of this example scene, consisting of the coordinates and type of each object present. The constraint inference step does not involve a PDDL problem description. Once constraints have been generated, they are used to plan on specific problem instances. At this stage, the specific PDDL problem description is accessed by the task planner.
We use the PDDL standard for symbolic task planning. For a given domain, a PDDL domain specifies the allowed types, predicates and actions. Actions are defined by their preconditions, expressed as logical formulas over predicates, and their effects, which specify how the truth values of a subset of predicates are updated in the next state. An instance of a problem in a particular domain defines a set of objects, their initial state, and the goal conditions. States are represented as a set of predicates which true. We adopt the closed-world assumption, in predicates not specified to be true are implicitly assumed to be false. The problem of task planning is to find a sequence of actions which achieves the goal from the initial state.
@@ IV Methodology
Figure 3: VIZ-COAST’s Visual Reasoning Module infers constraints through a 4-step prompting procedure. First, it is asked to provide an interpretation of the scene based on an image of the example problem instance and the PDDL domain description. Then, it is asked to articulate in natural language the necessary constraints. After that, the VLM is asked to formally encode the constraints it identified in Python using Z3’s API, referencing a structural example. Finally, it proofreads the script to ensure syntactic compliance.
@@
  • SMT-based Task planner: We use an Satisfiability Module Theories (SMT) solver—Z3 [8], in particular—that uses a SATPlan encoding [6, 7] to enable flexible constraint specification while introducing minimal search overhead. This is similar to the task planning framework used in IDTMP [5] and CaStL [33].
  • Visual Reasoning Module (VRM): Before any planning occurs, the VRM analyzes an image of the scene paired with the PDDL domain specification to foresee downward refinement failures and infer general spatial constraints to prevent them. These are specified as constraints on the SMT solver.
  • Stream-based Grounding: We ground our task plans with the stream-based planning used PDDLStream [4] and COAST [9].
We first briefly introduce the task planning module, then we describe the visual reasoning module in detail, and how it interacts with the downstream task and motion planning components. Our primary contribution lies in the VRM’s ability to foresee and prevent downward refinement failures.
@@ IV-1 Task planning
VIZ-COAST employs a general-purpose task planner based on the Z3 Satisfiability Modulo Theories (SMT) solver [8]. We use the Madagascar [7] encoding for SATPlan [6]. This allows us to define declarative SMT constraints on allowable actions and predicates through Z3’s Python API, taking the form of arbitrary expressions of first-order logic. Other popular planners such as Fast Downward [21] do not have a flexible interface for constraint specification, and require editing the PDDL files directly, introducing additional unnecessary complexity, especially due to the translation step to SAS+ [36] required.

### v2

@@ Abstract
In task and motion planning, high-level task planning is done over an abstraction of the world to enable efficient search in long-horizon robotics problems. However, the feasibility of these task-level plans relies on the downward refinability of the abstraction into continuous motion. When a domain’s refinability is poor, task-level plans that appear valid may ultimately fail during motion planning, requiring replanning and resulting in slower overall performance. Prior works mitigate this by encoding refinement issues as constraints to prune infeasible task plans. However, these approaches only add constraints upon refinement failure, expending significant search effort on infeasible branches. We propose VIZ-COAST, a method of leveraging the common-sense spatial reasoning of large pretrained Vision-Language Models to identify issues with downward refinement a priori, bypassing the need to fix these failures during planning. Experiments on three challenging TAMP domains show that our approach is able to extract plausible constraints from images and domain descriptions, drastically reducing planning times and, in some cases, eliminating downward refinement failures altogether, generalizing to a diverse range of instances from the broader domain.
@@ I Introduction
Task and Motion Planning (TAMP) [garrett2021integrated] offers a framework for solving long-horizon robotic problems that couples symbolic decision-making with geometric feasibility, such as clearing a cluttered table or navigating between rooms where a key must first be found to open a door. These problems are compelling not only because tasks require multiple steps, but also because they demand reasoning across different levels of abstraction. The robot must determine what actions to take and how to execute them. This feature makes TAMP inherently challenging as it requires reconciling discrete task structures with high-dimensional geometric constraints.
@@
The two levels of abstraction, task and motion, enable efficient planning in long-horizon problems. At the task level, actions are represented symbolically, such as “pick,” “place,” or “navigate,” with each action specified by preconditions and effects on a set of symbolic predicates. This abstraction allows the planner to efficiently reason about possible sequences of actions without searching in continuous geometric space. However, the core difficulty is that these abstractions often cannot express critical geometric information, violating the downward refinement property [bacchus1991downward, bacchus1994downward]: a valid task plan may not necessarily correspond to a sequence of feasible motion plans. For instance, a symbolic plan may prescribe grasping an object, but that object may be obstructed by another object, preventing any collision-free grasp.
@@
Figure 1: VIZ-COAST uses VLMs and SMT-based task planning to constrain TAMP problems. Our constraints generalize to unseen problem instances within a broader domain. VIZ-COAST introduces minimal search overhead, enabling more efficient planning than prior state-of-the-art.
@@
PDDLStream [garrett_pddlstream_2020] is a TAMP framework which circumvents downward refinement failures by employing a sample-first approach, in which feasible continuous action parameterizations, such as grasp poses and trajectories, are sampled prior to planning via streams, allowing task planning to be done with geometric feasibility pre-certified. However, PDDLStream expends significant effort sampling action groundings which may not be used in the final plan. This can be very inefficient, as the number of possible action groundings grows exponentially with the number of objects.
@@
Alternatively, many TAMP methods take a plan-first approach, creating a complete task plan before attempting downward refinement. Intuitively, it is far more effective to prune the search space at the discrete task level than the continuous motion level, especially in long-horizon tasks. When refinement fails, plan-first methods translate motion failures to task-level constraints before replanning. For example, upon failing to find a feasible grasp pose due to an obstructing object, a constraint may be added to enforce that the offending object must not be in that position when attempting a grasp. IDTMP [dantam2018incremental] implements this by formulating planning with SATPlan [kautz2006satplan, rintanen2014madagascar] through Z3 [de2008z3], a Satisfiability Modulo Theories (SMT) solver, incrementally adding constraint clauses upon failure. COAST [vu2024coast], a plan-first successor to PDDLStream, adds auxiliary predicates and entities to directly encode constraints in the PDDL description. All these existing constraint-based approaches ultimately require expending additional search effort to encounter downward-refinement failures and discover the necessary constraints, often replanning several times before finding a plan that can be successfully ground to motion. Our goal is to avoid expending this effort by adding the necessary constraints before planning even begins.
@@
Recently, large pretrained vision language models (VLMs) have demonstrated common-sense geometric and physical reasoning capabilities [gao2024physically, chen2024spatialvlm], allowing them to generalize zero-shot to a wide range of settings. For this reason, they have been employed for various facets of TAMP, including subgoal generation [yang_guiding_2025], encoding open-world preferences [kumar_open-world_2025], and correcting errors [duan2024aha].
@@
In this work, we propose Visual Insight for Z3 [de2008z3]-based Constraints and Streams (VIZ-COAST), a novel approach to TAMP which leverages VLMs to foresee potential downward-refinement failures and add constraints to avoid them before planning begins, eliminating the need to encounter these failures during planning and expend effort replanning. In challenging domains where initial constraints do not account for all failure modes, VIZ-COAST analyzes multiple problems to revise constraints using feedback from downstream planning in a closed-loop manner.
@@
Through experiments on three challenging TAMP domains, we demonstrate that VIZ-COAST significantly decreases planning time by reducing or even eliminating replanning caused by downward refinement failures. We also show that our constraints consistently generalize to unseen problem instances within a broader domain.
@@ II Background and Related Work
We begin with a general overview of modern task and motion planning and its main challenges in Sec.˜II-A. We then review prior approaches in Sec.˜II-B that accelerate TAMP by constraining the search space, highlighting the need for a generalizable method capable of pruning before planning. Lastly, in Sec.˜II-C, we explain how VLMs’ spatial reasoning capabilities make such a method possible and clarify how our approach differs from prior work.
@@ II-A Task and Motion Planning
Task and motion planning (TAMP) decomposes continuous long-horizon motion planning into search in both discrete (task-level) and continuous (motion-level) spaces [garrett2021integrated]. The task level planning is done over an abstraction of the robot’s capabilities, and produces a sequence of symbolic actions to transition from a given start state to a desired goal condition (e.g., STRIPS [fikes1971strips]). In this work, we formalize task planning using Planning Domain Definition Language (PDDL) [mcdermott1998pddl, fox2003pddl2] due to its expressive power and widespread adoption.
A core challenge in TAMP is downward-refinement failure [bacchus1991downward, bacchus1994downward], which occurs when a symbolic task plan becomes infeasible at the motion level. PDDLStream [garrett_pddlstream_2020] addresses this issue by taking a sample-first approach, performing continuous sampling before task planning via declarative streams, allowing earlier reasoning about geometric feasibility. Other works [kaelbling_hierarchical_2011, srivastava2014combined, dantam2013motion, vu2024coast] take a plan-first approach, designing algorithms to add constraints upon encountering failures which inform task-level replanning to prevent revisiting infeasible options. In the next section, we expand on the role of constraints and other methods of reducing the search space in TAMP.
@@ II-B Constraints in TAMP
The purpose of constraints in both IDTMP and COAST is to prune the task-level search space to avoid repeating failures, though they differ in implementation. COAST [vu2024coast] is a plan-first adaptation of PDDLStream which encodes constraints by modifying the PDDL description, adding auxiliary predicates and entities to block actions under specific conditions. While this approach allows COAST to be planner-agnostic, encoding constraints directly in PDDL limits the constraints that may be expressed, and the auxiliary constructions introduce significant search overhead. PDDL 3.0 [Gerevini2005PDDLConstraints] introduced constraints on valid states, but is poorly supported by existing planners (e.g., Fast Downward [helmert2006fast]) and cannot be reliably generated by LLMs due to limited training data. In contrast, we take the approach used by IDTMP [dantam2018incremental], and solve for valid task plans through SATPlan [kautz2006satplan, rintanen2014madagascar] encoded in a Satisfiability Modulo Theories (SMT) solver (in particular, Z3 [de2008z3]). This formulation enables constraints to be directly encoded in the SMT solver through simple Python programming, enabling efficient and flexible constraint specification. While heuristic search approaches typically exhibit faster planning times [taitler20242023], SATPlan encodings offer far greater specification flexibility and can avoid the costly translation overhead in FastDownward [helmert2006fast] for many problems.
Although existing methods [vu2024coast, dantam2018incremental] are able to overcome downward refinement issues, they discover constraints only by encountering motion grounding failures during execution. This reactive process is often repeated for several iterations before uncovering the necessary constraints to solve a particular problem instance, and cannot generalize to unseen instances.
Alternative approaches have explored predicting action feasibility to avoid refinement failures. wells2019learning learn feasibility constraints for tabletop manipulation through domain-specific training, while xu2022accelerating employ neural feasibility checks to identify infeasible actions before motion planning. driess2020deep use deep visual heuristics to assess feasibility in mixed-integer manipulation planning, and [20-driess-RSS] leverage visual reasoning to predict valid actions for TAMP. Although promising, these methods require extensive domain-specific training data and may not generalize well to new environments or object configurations. In contrast, recent advances in LLMs and VLMs enable the use of pre-trained common-sense priors to reduce unnecessary exploration.
@@ II-C Vision Language Models for TAMP
LLMs have demonstrated that they are suitable for high-level reasoning [aghzal2025survey]. VLMs further integrate visual and textual inputs to enable scene understanding [shek2024lancar], spatial reasoning [chen2024spatialvlm], and grounded action execution [liu2024vision, duan2024aha]. These capabilities make VLMs especially promising for addressing the geometric reasoning challenges in TAMP.
Several works have integrated LLMs and VLMs into TAMP pipelines for various purposes. zhang2023grounding use VLMs to ground task planners by validating symbolic predicates and providing incremental feedback. hu2023look uses a VLM to generate physically feasible actions. ReplanVLM [mei2024replanvlm] employs VLMs for plan generation with replanning upon grounding failures. CaStL [guo_castl_2025] translates natural language task descriptions into PDDL with constraints reflecting user preferences. OWL-TAMP [kumar_open-world_2025] encodes open-world preferences as constraints on symbolic and geometric states. Both CaStL and OWL-TAMP operate at the problem specification level, aiming to capture preferences rather than prevent downward refinement failures. VLM-TAMP [yang_guiding_2025] uses VLMs to generate intermediate subgoals for long-horizon planning. While this accelerates task-level plan search, it relies on the assumption of downward refinability. CLIMB [byrnes_climb_2025] incrementally updates the domain description upon failure with an LLM to continually improve task performance; however, it is restricted to constraints expressible in PDDL, and does not make use of visual input.
@@ III Problem Formulation
Our work investigates the use of VLMs to accelerate TAMP. In particular, our method is capable of producing constraints which generalize to a range of problem instances within a broader domain.
When inferring constraints for a particular domain, we assume access to a model of the robot and its environment, as well as a symbolic Planning Domain Definition Language (PDDL) [mcdermott1998pddl] domain description. Additionally, we assume access to some example scene within the domain, consisting of several components: (1) a symbolic PDDL problem description; (2) the geometric coordinates and dimensions of each object present; (3) an image of the initial state of the scene which clearly shows all objects. Lastly, we assume the set of possible unique object instances is finite and known. Once constraints have been generated, they are used to plan on subsequent problem instances.
We use the PDDL standard for symbolic task planning. For a given domain, the PDDL domain description specifies the allowed types, predicates and actions. Actions are defined by their preconditions, expressed as logical formulas over predicates, and their effects, which specify how the truth values of a subset of predicates are updated in the next state. An instance of a problem in a particular domain defines a set of objects, their initial state, and the goal conditions. States are represented as the set of predicates which are true. We adopt the closed-world assumption, in which predicates not specified to be true are implicitly assumed to be false. The problem of task planning is to find a sequence of actions which achieves the goal from the initial state.
@@ IV Methodology
Figure 2: The VIZ-COAST architecture. Our Visual Reasoning Module takes as input an example scene, consisting of an image and the geometric state, the PDDL domain and problem description, and the grasp information in natural language. It produces a Python file encoding constraints through a high-level API to the SMT-based task planner. When presented with a new problem instance, the task planner takes these constraints as input, in addition to the PDDL domain and new problem descriptions. The planner applies the constraints to block geometrically infeasible task plans. The high-level plan produced is then grounded by a streams-based motion planner to produce an executable continuous motion plan.
@@
  • SMT-based Task planner: We use a Satisfiability Modulo Theories (SMT) solver—Z3 [de2008z3], in particular—that uses a SATPlan encoding [kautz2006satplan, rintanen2014madagascar] to enable flexible constraint specification while introducing minimal search overhead. This is similar to the task planning framework used in IDTMP [dantam2018incremental] and CaStL [guo_castl_2025].
  • Visual Reasoning Module (VRM): Before planning begins, the VRM uses a VLM to analyze an example scene to foresee downward refinement failures and infer general constraints to prevent them, encoded as constraints on the SMT solver.
  • Stream-based Grounding: We ground task plans to motion with the streams framework used in PDDLStream [garrett_pddlstream_2020] and COAST [vu2024coast].
We first briefly introduce the task planning module, then we describe the VRM in detail, and how it interacts with the downstream task and motion planning components. The full architecture is shown in Fig.˜2. Our primary contribution lies in the VRM’s ability to foresee and prevent downward refinement failures.
@@ IV-1 Task planning
VIZ-COAST employs a general-purpose task planner based on the Z3 SMT solver [de2008z3]. We use the Madagascar [rintanen2014madagascar] encoding for SATPlan [kautz2006satplan]. This allows us to define declarative SMT constraints on allowable actions and predicates through Z3’s Python API, taking the form of arbitrary expressions of first-order logic. Other popular planners such as Fast Downward [helmert2006fast] do not have a flexible interface for constraint specification, and require editing the PDDL files directly, introducing additional unnecessary complexity, especially due to the translation step to SAS+ [jonsson1998state] required.